### v1

# Short-time simulation of quantum dynamics by Pauli measurements

Abstract
Simulating the dynamics of complex quantum systems is a central application of quantum devices. Here, we propose leveraging the power of measurements to simulate short-time quantum dynamics of physically prepared quantum states in classical post-processing using a truncated Taylor series approach. While limited to short simulation times, our hybrid quantum-classical method is equipped with rigorous error bounds. It is extendable to estimate low-order Taylor approximations of smooth, time-dependent functions of tractable linear combinations of measurable operators. These insights can be made use of in the context of Hamiltonian learning and device verification, short-time imaginary time evolution, or the application of intractable operations to sub-universal quantum simulators in classical post-processing.
I Introduction
The study of the dynamics of quantum systems is a highly active field of research and predicting dynamical properties constitutes one of the most anticipated tasks to be tackled by quantum devices. Recent developments have seen an impressive rise in experimentally available qubit numbers and quantum operations in actual physical systems, be it for universal quantum devices [1] or sub-universal quantum simulators [2]. However, full-fledged quantum simulation is still prohibitively expensive, and purely classical methods are limited since a universal, efficient classical algorithm for quantum simulation would imply that all quantum algorithms could be efficiently simulated classically (${\rm BPP}={\rm BQP}$). This is because quantum dynamics generated by local Hamiltonians is ${\rm BQP}$ complete and hence as computationally powerful as a full quantum computer when formalized as a decision problem [3, 4, 5]. To make the best use of the currently available resources, recent developments have thus focused on finding resource-efficient algorithms, be it through hybrid quantum-classical approaches outsourcing parts of the computation to classical hardware, designing algorithms tailored to specific hardware and simulation tasks, or improving classical methods. In this delicate and important regime, trade-offs between the required circuit depth and the number of required circuit evaluations are also of key interest, since the former is highly limited in current, noisy quantum devices.
Following a second line of thought, in recent years, the notion of randomized measurements has gained traction in a variety of specific flavors, following a “measure now, compute functions of interest later” mindset [6, 7, 8, 9] that is essential to the efficient functioning of most hybrid quantum-classical algorithms that can be executed on near-term hardware. Besides estimating the energy of some trial states necessary for the variational quantum eigensolver and most other variational quantum algorithms [10, 11, 12], such methods can also be used to estimate expectation values of observables with small Frobenius norms, such as overlaps with pure states [6, 13] or estimating energy gaps with no additional resources beyond time evolution and measurements [14]. These applications thus impressively showcase the fundamental power of quantum measurements and classical post-processing and highlight the need for dedicated readout schemes such as grouping strategies [15], graph optimization techniques [16] or fully randomized or de-randomized schemes [6, 7, 17, 18]. After all, the power of quantum measurement goes hand in hand with efficient measurement protocols.
Randomness has found even more applications in reducing resource costs of near-term and early fault-tolerant quantum devices as it can help reduce errors that arise for product formulas [19, 20, 21, 22, 23]. Furthermore, in the context of algorithms that use what are called linear combinations of unitaries (LCU) [24, 25, 26, 27, 28, 23], randomness can help circumvent the need for block encodings [23, 29, 30]. For certain settings, these LCUs can be constructed purely in post-processing [31] without actually requiring any auxiliary qubits.
While these methods approach the task of making quantum simulation more feasible from the side of improving quantum algorithms or trading resources to accommodate noisy and early fault-tolerant quantum devices, there has also been a lot of recent progress toward improved purely classical methods simulating quantum dynamics, using tensor network methods [32, 33, 34], truncated Pauli paths [35] or cluster expansions [36]. Cluster expansion methods are well-suited for the classical simulation of short-time quantum dynamics for product states evolving under a few-body Hamiltonian. They are reminiscent of the quantum algorithm simulating Hamiltonian dynamics with truncated Taylor series [27], another LCU approach, and use nested commutators to reduce computational costs. These conceptual similarities raise the crucial question of whether such methods can be used in a hybrid quantum-classical setting to help enhance current quantum devices.
This work demonstrates how the power of randomized Pauli measurements can be leveraged to estimate short-time dynamics of arbitrary quantum states in classical post-processing. Moreover, this framework readily extends beyond quantum dynamics and applies to any low-order Taylor approximation of spectral functions of linear combinations of measurable operators, not just Pauli operators. As such, we anticipate further practical applications in notions of Hamiltonian learning that are based on short-time dynamics or in device verification, in short-time imaginary time evolution, or for introducing operations intractable to a given sub-universal quantum simulator in classical post-processing. These applications will be sketched here.
II Methods
II.1 Problem setting
The precise problem we address in this work is the following: Consider an $n$-qubit system whose Hamiltonian $H$ be specified as the linear combination of $L$ Pauli operators $H_{\ell}$ acting on an $n$-qubit system, that is,
$H=\sum_{\ell=1}^{L}\alpha_{\ell}H_{\ell},\quad\text{with }\quad\lambda=\sum_{\ell=1}^{L}|\alpha_{\ell}|,$ (1)
where the number of non-trivial tensor factors per Pauli operator, also denoted as their Pauli weight, is upper bounded by $w$. Furthermore, let $O$ be some (linear) observable. Then, the goal is to estimate the time dynamics $\mathrm{tr}\!\left(\mathrm{e}^{\mathrm{i}Ht}O\mathrm{e}^{-\mathrm{i}Ht}\rho\right)$ for an arbitrary, potentially mixed, initial state $\rho$ and (short-)time $t>0$.
In a more general sense, we are interested in estimating quantities of the form $\mathrm{tr}\!\left(F(H,t)\rho\right)$, where $(H,t)\mapsto F(H,t)$ can be any function depending on a linear combination of Pauli operators $H$ and variable $t$ whose truncated Taylor series with respect to $t$ can be bounded. To avoid a convoluted presentation, we will focus on the specific application of estimating time dynamics and mention how the procedure would differ in a more general setting.
II.2 Expanding time dynamics with a Taylor series
Using the approach of employing a truncated Taylor series for Hamiltonian simulation proposed in Ref. [27], we can approximate such functions by implementing their Taylor series truncated at $K$-th order, i.e.,
$U(t)$ $\coloneqq\mathrm{e}^{-\mathrm{i}Ht}\approx\sum_{k=0}^{K}\frac{(-\mathrm{i}Ht)^{k}}{k!}$
$\approx\sum_{k=0}^{K}\sum_{\ell_{1},\ldots,\ell_{k}=1}^{L}\underbrace{\frac{(-\mathrm{i}t)^{k}}{k!}\alpha_{\ell_{1}}\cdots\alpha_{\ell_{k}}}_{\beta_{j}}\underbrace{H_{\ell_{1}}\cdots H_{\ell_{k}}}_{P_{j}}$ (2)
$\widetilde{U}(t)$ $=\sum_{j=1}^{m}\beta_{j}P_{j}$ (3)
which itself is the sum of $m\leq\sum_{k=0}^{K}L^{k}=\left(L^{K+1}-1\right)/(L-1)=\mathcal{O}\left(L^{K}\right)$ $n$-qubit Pauli operators $P_{j}$ corresponding to some $H_{\ell_{1}}\cdots H_{\ell_{k}}$ (whose weight is upper bounded by $\min\{wK,n\}$) and associated pre-factors $\beta_{j}$. As long as $\lVert H\rVert t\leq\lambda t\leq 1$, e.g., in the case of short-times, the accuracy of this approximation is upper-bounded by $\epsilon$ when choosing $K=O(\log(1/\epsilon))$. Appendix A thoroughly discusses and analyzes the truncation error of a Taylor approximation of the time evolution operator, recovering previous results of Ref. [27]. However, this is not the case anymore for longer simulation times. One way to circumvent this issue is to slice the time evolution into $r\geq\lVert H\rVert t$ segments $U(t)=U(t/r)^{r}$, each with accuracy $\epsilon/r$, increasing the number of terms in Eq. (3) by a factor of $r$ as one would for product formula approaches.
Note that the time-evolved observable $F(H,t)=\mathrm{e}^{\mathrm{i}Ht}O\mathrm{e}^{-\mathrm{i}Ht}=O(t)$ features two unitary evolutions. We can approximate both of them separately and obtain
$\widetilde{O}(t)$ $=\widetilde{U}(-t)O\widetilde{U}(t)=\sum_{j,j^{\prime}=1}^{m}\underbrace{\beta_{j}\beta_{j^{\prime}}}_{\gamma_{i}}\underbrace{P_{j}OP_{j^{\prime}}}_{Q_{i}}$ (4)
$=\sum_{i=1}^{m^{2}}\gamma_{i}Q_{i},$ (5)
which is now a sum of at most $m^{2}$ different $n$-qubit Pauli operators of the form $Q_{i}=P_{j}OP_{j^{\prime}}H_{\ell_{1}}\cdots H_{\ell_{k}}OH_{\ell_{1}}\cdots H_{\ell_{k^{\prime}}}$. A priori, the coefficients $\gamma_{i}=\beta_{j}\beta_{j^{\prime}}$ can be complex-valued, but the underlying Hermitian structure of the time-evolved observable (i.e. symmetry under taking adjoints) ensures that complex-valued coefficients must cancel out and the total sum ranges over $m_{\mathrm{tot}}\leq m^{2}$ terms with real-valued coefficients $\gamma_{i}\in\mathbb{R}$.
Estimating $\mathrm{tr}\!\left(O(t)\rho\right)$ is thus well approximated by a real-valued, linear combination of the expectation values of Pauli operators $Q_{i}$, i.e.,
$\mathrm{tr}\!\left(O(t)\rho\right)\approx\mathrm{tr}\!\left(\widetilde{O}(t)\rho\right)=\sum_{i=1}^{m_{\mathrm{tot}}}\gamma_{i}\mathrm{tr}\!\left(Q_{i}\rho\right),$ (6)
where $m_{\mathrm{tot}}$, the total number of Pauli operators to be estimated, can be significantly smaller than the upper bound of $m^{2}$ from Eq. (5).
For $K=O(\log(r/\epsilon))$ and $r=\lVert H\rVert t$, the number of Pauli expectation values to be estimated is upper bounded by
$m_{\mathrm{tot}}\leq\left(\sum_{k=0}^{K}L^{k}\right)^{2r}\!\!=\mathcal{O}(L^{2Kr})=\mathcal{O}\left(\left(\frac{\lVert H\rVert tL}{\epsilon}\right)^{2\lVert H\rVert t}\right),$ (7)
where $\|H\|$ denotes the operator (or spectral) norm of the Hamiltonian $H$. The maximum Pauli weight over all $Q_{i}$s is given by
$w_{\mathrm{max}}=\min(\mathcal{O}\left(2ww_{O}\lVert H\rVert t\log(\lVert H\rVert t)),n\right).$ (8)
Estimating time dynamics can, therefore, effectively be reduced to a pure measurement task – relevant in many meaningful physical platforms – albeit with potentially many required measurements. Since the number of expectation values of Pauli operators to be estimated grows exponentially in the simulation time, this method is practically limited to short simulation times.
[…]
It is important to note that since the Taylor expansions rely on powers of the Hamiltonian and thus powers of sums of Pauli operators, not all will be pairwise different. This is due to their group structure and the fact that Pauli operators square to the identity. Consequently, the effective number of terms in this sum can be significantly smaller than the provided upper bound. However, an exponential scaling in the chosen cutoff $K$ and number of segments $r$ cannot be fully avoided.
For time-evolved observables $O(t)$ in particular, using the fact that the time evolution in the Heisenberg picture for time-independent Hamiltonians and observables results in well-behaved nested commutators, we find that Eq. (25) can be rewritten as
$\mathrm{e}^{\mathrm{i}Ht}O\mathrm{e}^{-\mathrm{i}Ht}=O+\mathrm{i}t[H,O]+\frac{(\mathrm{i}t)^{2}}{2!}[H,[H,O]]+\ldots,$ (9)
which is reminiscent of cluster expansion techniques [36] using nested commutators to decide which terms to estimate in a classical algorithm on product states, and stressing that we are leveraging the group structure of the Pauli operators. The advantage of such a reformulation is that we no longer have to calculate all combinations of Pauli operators to obtain the $Q_{i}$ in Eq. (5), many of which will cancel, but can use the underlying structure to calculate only non-vanishing terms. Using the knowledge that $\mathrm{tr}\!\left(O(t)\rho\right)$ has to be real-valued, we can further eliminate all odd powers of $t$. The connection to well-understood cluster expansion techniques provides a straightforward path to enhancing such techniques with quantum resources. Instead of performing such an expansion and estimating the individual Pauli expectation values classically for product states as in Ref. [36], we can measure them on arbitrary quantum states using existing expansions. Since nested commutators are a reformulation of the Taylor series of $O(t)$, the same error bounds as in Appendix B apply. Central to cluster expansion techniques and extensions thereof, thorough discussions can be found, e.g., in Refs. [36, 38]. Knowledge of the structure of the function of interest can thus help reduce the amount of classical post-processing.
Instead of concatenating two Taylor approximations of $U(t)$ for $t>0$ into an approximation of $O(t)$, as we have presented above, we can also directly Taylor expand the target function, as one would do for more generic functions $F(H,t)$. How such a direct expansion would present itself and differences in accuracy and the number of terms are discussed in Appendix B.
II.3 Error bounds
An approximation of $\mathrm{tr}\!\left(F(H,t)\rho\right)$ via a truncated Taylor series followed by an estimation of the resulting linear combination of expectation values of Pauli operators, which is accompanied by two types of errors: A systematic error due to the truncation of the Taylor series and a sampling error obtained during the estimation of the Pauli expectation values. Both of these error types can be rigorously bounded.
II.3.1 Systematic errors
We can upper-bound the operator norm difference between the truncated Taylor series and target matrix function $F(H,t)$ by employing Taylor’s remainder theorem as showcased in Appendix A for a direct Taylor expansion of the particular matrix-valued function $F(H,t)=O(t)$. When stitching together multiple approximations as in Eq. (4) above, where we have constructed an approximation of $O(t)$ using truncated Taylor series for $U(t)$, we can also combine the individual approximation errors to bound the total error when approximating the target function. As in Ref. [23, Lemma 1], given an approximation $\widetilde{U}$ of $U$ with $\lVert U-\widetilde{U}\rVert\leq\epsilon$, we can redefine $\widetilde{U}=U+U_{\epsilon}$ with $\lVert U_{\epsilon}\rVert\leq\epsilon$. We then find
$\left\lVert U^{\dagger}OU-\widetilde{U}^{\dagger}O\widetilde{U}\right\rVert$ $=\left\lVert U_{\epsilon}^{\dagger}OU+U^{\dagger}OU_{\epsilon}+U_{\epsilon}^{\dagger}OU_{\epsilon}\right\rVert$ (10)
$\leq 2\lVert U\rVert\lVert U_{\epsilon}\rVert\lVert O\rVert+\lVert U_{\epsilon}\rVert^{2}\lVert U\rVert$
$\leq(2\epsilon+\epsilon^{2})\lVert O\rVert$
$\leq 3\epsilon\lVert O\rVert,$
providing a rigorous error bound for the systematic error of the approximation $\widetilde{O}(t)$ defined in Eq. (4) to the actual time dynamics. Using techniques from Ref. [39], the error bound could be further tightened using nested commutators.
These bounds constitute only the systematic error originating in the truncation of the Taylor series and not the error obtained in implementing the linear combinations of Pauli operators, which is method-dependent and would have to be added to bound the actual error picked up during implementation.
II.3.2 Sampling error
Using state-of-the-art methods, the linear combination of unitaries of Eq. (5) could be implemented deterministically using block encodings in the spirit of Refs. [37, 27] or in a randomized fashion using the frameworks of Refs. [30, 23]. However, since we are dealing with sums of Pauli operators, simply performing Pauli basis measurements as outlined in Eq. (6) suffices and no further quantum circuitry is required.
Thus, estimating $\mathrm{tr}\!\left(\mathrm{e}^{\mathrm{i}Ht}O\mathrm{e}^{-\mathrm{i}Ht}\rho\right)$ is equal to estimating the expectation values of the sum of Pauli operators given by the nested commutator in Eq. (9) for some initial state $\rho$. Therefore, besides preparing the input state $\rho$, the main cost of this algorithm lies in the number of measurements required to obtain sufficient statistical data for all terms of the truncated Taylor series.
This sampling error can be approached with two different mindsets: we can view these methods as the primary objective of an experiment or as secondary to another task. That is, we either care only about $\mathrm{tr}\!\left(F(H,t)\rho\right)$ or are performing another experiment, such as a variational algorithm estimating $\mathrm{tr}\!\left(H\rho(\bm{\theta})\right)$ for some trial state $\rho(\bm{\theta})$, using some readout scheme that produces classical data from which one could, albeit sub-optimally, also infer information about $\mathrm{tr}\!\left(F(H,t)\rho\right)$.
Direct estimation
If this method is the primary objective, we want to reduce the sampling error as much as possible while also being practical, i.e., reducing the classical pre-processing cost. To this end, we propose using a randomized measurement strategy, although any measurement scheme is viable. The advantage of a randomized scheme is that each sample taken from the quantum devices will then be an unbiased estimator of the whole sum with corresponding convergence properties. To simplify notation, we denote the Taylor expansion as a linear combination of weighted Pauli operators, pulling the sign of each pre-factor into their corresponding Pauli, i.e.,
$\widetilde{O}(t)$ $=\sum_{i=1}^{m_{\mathrm{tot}}}\gamma_{i}Q_{i}=\sum_{i=1}^{m_{\mathrm{tot}}}\lvert\gamma_{i}\rvert(\operatorname{sign}(\gamma_{i})Q_{i})$ (11)
$=\lVert\gamma\rVert_{1}\sum_{i=1}^{m_{\mathrm{tot}}}\underbrace{\frac{\lvert\gamma_{i}\rvert}{\lVert\gamma_{i}\rVert_{1}}}_{p_{i}}\underbrace{(\operatorname{sign}(\gamma_{i})Q_{i})}_{P_{i}},$
where we further normalize the pre-factors $\gamma_{j}$ such that we arrive at a probability distribution $\{p_{i},P_{i}\}$. Sampling which Pauli expectation value $\mathrm{tr}\!\left(P_{i}\rho\right)$ to measure in each circuit evaluation from this distribution results in an unbiased estimator of $\widetilde{O}(t)$. Using Hoeffding’s inequality, we then find that
$N\geq\frac{2\lVert\gamma\rVert_{1}^{2}\ln(2/\delta)}{\epsilon^{2}}$ (12)
samples suffice to estimate the linear combination of Pauli operators up to an error of $\epsilon$ with probability at least $1-\delta$.
The estimation can be improved using structural knowledge of the target function, e.g., by incorporating the nested commutator structure and quantum resources saved by immediately returning zero for all complex-valued $\gamma_{j}$ when considering $O(t)$. Alternatively, grouping strategies [15], graph optimization techniques [16], fully randomized measurements [6, 7] or other dedicated readout schemes could be used with their corresponding sampling error bounds.
Using existing measurement data
Especially quantum algorithms using randomized classical shadows [6, 7, 8, 9] are of additional interest. Since these strategies by default aim to estimate a set of $m$ $n$-qubit Pauli operators $P_{i}$, they constitute the prototypical “measure now, compute functions of interest later” mindset. Thus, even if the main objective of a quantum algorithm differs from those discussed here, one still has the measurement data to reuse and can formulate a sampling error bound given the number of samples.
The performances of these methods are shown mostly heuristically with an analytical upper bound on the number of measurements given by Ref. [40, Theorem 3]. It essentially states that for an accuracy $\epsilon\in(0,1)$ and failure probability $\delta$, it suffices to (sequentially) use
$N=\frac{2}{\epsilon^{2}(1-\epsilon)}3^{w}\log\frac{3L}{\delta}$ (13)
copies of $\rho$ to approximate each $\mathrm{tr}\!\left(P_{i}\rho\right)$ up to (additive) accuracy $\epsilon$, where $w$ denotes the maximum Pauli weight.
For the presented algorithm these results lead to the following conclusion: controlling the error of each $Q_{i}$ in Eq. (6) to $\epsilon/\lVert\gamma_{i}\rVert_{1}$ and thus bounding the overall sampling error of $\widetilde{O}(t)$ by $\epsilon$ requires
$N=\mathcal{O}\left(3^{w_{\mathrm{max}}}\frac{\lVert\gamma\rVert_{1}^{2}\log(m_{\mathrm{tot}}/\delta)}{\epsilon^{2}}\right)$ (14)
copies of $\rho$. While the actual performance could be better by orders of magnitude, these methods usually perform best for many low-weight or few global terms in contrast to the Taylor series which outputs many terms with intermediate to large Pauli weight. Having obtained measurement data for another, primary, task, we can reuse the measurement data to secondarily estimate matrix functions $(H,t)\mapsto F(H,t)$ up to an explicitly determinable error.
III Concrete example and reducing terms
To gain better intuition of how the expansion of time dynamics into a linear combination of Pauli operators of Eq. (5) might look like for a specific Hamiltonian, let us consider a Heisenberg Hamiltonian with equally weighted interactions, Take for example a one-dimensional Heisenberg Hamiltonian with equally weighted interactions, i.e.,
$H=\sum_{i=1}^{n-1}J\left(X_{i}X_{i+1}+Y_{i}Y_{i+1}+Z_{i}Z_{i+1}\right),$ (15)
which has three Pauli operators per site and thus a total number of $L=3(n-1)$ terms.
Naively, we have upper-bounded the number of terms present in the expansion of the $K$-th order truncated Taylor series of the time evolution operator as $m_{\mathrm{tot}}=\sum_{k=0}^{K}L^{k}=(L^{K+1}-1)/(L-1)$. However, since the Pauli operators square to the identity and also form a group, many terms will collapse to the identity matrix or reproduce already existing terms as early as in the second order. Explicitly, for
$H^{2}=$ $\left(\sum_{i=1}^{n-1}\underset{h_{i}}{\underbrace{J\left(X_{i}X_{i+1}+Y_{i}Y_{i+1}+Z_{i}Z_{i+1}\right)}}\right)^{2}$
$=$ $\sum_{i,j=1}^{n-1}h_{i}h_{j}=\sum_{i=1}^{n-1}h_{i}^{2}+\sum_{i<j}(h_{i}h_{j}+h_{j}h_{i})$
we find that
$\left(h_{i}\right)^{2}=3J^{2}\mathbb{I}-2J^{2}h_{i},$ (16)
which, instead of adding nine new Pauli operators, does not contribute to the growth of operators to be measured at all. Indeed, the identity term is already present in the zeroth order and the second term in the first order, leading to an overall reduction by $9(n-1)$ terms. Furthermore, terms of the form $X_{i}X_{i+1}X_{j}X_{j+1}$ for $i\neq j$ will appear twice in the second sum, further reducing the number of terms by $3n(n-1)$.
Thus, using algebraic properties of the Pauli operators, the total number of contributing terms for a Taylor series truncated after the second order is smaller than the worst-case bound of $(27(n-1)^{3}-1)/(3(n-1)-1)$ by $(n+9)(n-1)$, leading to $m_{\mathrm{tot}}=8n^{2}-23n+16$ terms. Note, however, that although these terms do not contribute to the number of Pauli operators, they still contribute their pre-factors which must be considered. These pre-factors $\gamma_{j}$ in Eq. (5) will then be sums of powers of $Jt$.
Further reductions in the number of contributing operators might be achieved by using a group different from the Pauli group as the basis for expanding the Hamiltonian and observable. Turning towards the multi-qubit Clifford group, we can simplify the Heisenberg Hamiltonian to
$H=\sum_{i=1}^{n-1}\frac{J}{2}\mathrm{SWAP}_{i,i+1}+(n-1)\frac{J}{2}\mathbb{I},$ (17)
using $\mathrm{SWAP}_{i,j}=(\mathbb{I}+X_{i}X_{j}+Y_{i}Y_{j}+Z_{i}Z_{j})/2$, and thereby reducing the number of terms in the Hamiltonian by a factor of three.
In general, finding appropriate decompositions is a hard task and thus not necessarily feasible, albeit a trick to be aware of when reducing circuit evaluations. Additionally, non-Hermitian operators, such as the Clifford operators, can not be directly measured and require additional circuitry such as the Hadamard test to perform measurements. As such, using different decompositions might be even more useful in the purely classical simulation regime.
IV Applications
Since the method proposed in Section II.2 relies on preparing an interesting quantum state and applying a function such as time evolution in classical post-processing, we view it as a hybrid quantum-classical algorithm showcasing the power of quantum measurements. Furthermore, any quantum computation will require measurements to output its result in a classically understandable language. Thus, we will always have access to at least a few measurements that could enhance the quantum computation in classical post-processing or as a standalone algorithm where we crank up the number of measurements to improve the accuracy of these methods.
IV.1 Short-time dynamics by measurement
The first practical application lies in quantum simulation. As a guiding example, we have already studied the use for estimating time dynamics of observables via measurements, reminiscent of classical cluster expansion techniques applied to arbitrary quantum states [36]. As such, any application of Cluster expansion, such as the Loschmidt echo $\mathrm{tr}\!\left(\rho\mathrm{e}^{-\mathrm{i}Ht}\right)$, is feasible. Overcoming the restriction to product states and moving towards a hybrid quantum-classical algorithm opens the door to new use cases. For near-term quantum devices, where time evolution algorithms are limited due to device noise but additional measurements might be realistic, we can imagine combining time evolution on the quantum device with additional short-time evolution in classical post-processing to enhance the quantum device capabilities.
To be more concrete, we can imagine a quantum circuit preparing the state $\rho(t_{1})=\mathrm{e}^{-\mathrm{i}Ht_{1}}\rho_{0}\mathrm{e}^{\mathrm{i}Ht_{1}}$ for the Néel state $\rho_{0}=\ket{0,1,0,1,\cdots}\!\bra{0,1,0,1,\cdots}$ | as initial quantum state and the Heisenberg Hamiltonian from Eq. (15). We can then estimate the expectation value of the particle imbalance $O=\sum_{i}(-1)^{i}Z_{i}$ not only for $\rho(t_{1})$ but also classically construct an approximation for $\mathrm{tr}\!\left(O\rho(t_{1}+t_{2})\right)$ for $t_{2}>0$ using the presented method.
As an intermediate regime between this hybrid algorithm and classical cluster expansion methods, we can also envision an extended cluster expansion that is not restricted to product states but to the superposition of $r$ product states, thereby increasing the computational cost, but at the same time remaining in the purely classical regime.
IV.2 Short-time Hamiltonian learning and device verification
Another example of practical short-time quantum simulating in post-processing is the task of device verification or Hamiltonian learning [41, 42, 43, 44]. Given a quantum device with an underlying, unknown Hamiltonian, we envision a protocol in which a well-known initial state is prepared and the system then evolves under its Hamiltonian for a short time before the state is measured. In classical post-processing, we can then evolve the system in negative time under a model Hamiltonian and check the statistics on certain observables to test the accuracy of the model Hamiltonian.
More concretely, we envision a protocol motivated by quantum process tomography and randomized benchmarking [45, 46, 47] in which we first prepare a basis state, e.g., $\rho=\ket{0}\!\bra{0}$ |. Then, we let the system naturally evolve the state as
$\rho(t)=\mathrm{e}^{-\mathrm{i}H_{\mathrm{sys}}t}\rho\mathrm{e}^{\mathrm{i}H_{\mathrm{sys}}t}.$ (18)
We can then gauge the accuracy of a guess $H_{\mathrm{guess}}$ of the system’s Hamiltonian $H_{\mathrm{sys}}$ with the presented method using $\mathrm{tr}\!\left(Z(-t)\rho(t)\right)$ for $Z(-t)=\mathrm{e}^{-\mathrm{i}H_{\mathrm{guess}}}Z\mathrm{e}^{-\mathrm{i}H_{\mathrm{guess}}}$, which would return zero if $H_{\mathrm{sys}}=H_{\mathrm{guess}}$. Repeating this procedure for different input states and observables then provides information about the accuracy of $H_{\mathrm{guess}}$
As such, it could be a quick device verification protocol restricted due to the short simulation times and less demanding on the hardware than extensive benchmarking protocols since it only relies on measurements and does not require additional quantum gates.
This method is, therefore, in the same spirit as Hamiltonian learning and device verification protocols using measurements of stationary states [41, 42, 43]. These use observables measured on states and their time evolution to construct an equation system from which the Pauli coefficients of the system’s Hamiltonian can be reconstructed. While these methods are tailored towards Hamiltonian learning, the proposed application of our Taylor series approach lies more in quick device verification.
IV.3 Short-time imaginary time evolution
Apart from real-time evolution, the nested commutator relations also hold for imaginary time evolution, as does the requirement on the truncation order (see, e.g., Ref. [48, Lemma 3.2]). Even though imaginary time evolution is mostly interesting for long simulation times due to its ability to drive a state toward its ground state, even a short simulation can be useful. In a setting where the initial state is prepared in a black-box fashion, e.g., a variational quantum algorithm or the output of some applied shortcut to adiabaticity, even short-time imaginary time evolution will push the prepared state closer toward the system’s ground state. As such, we can, with additional measurements, obtain measurement statistics that are closer to that of the true ground state in classical post-processing.
In more concrete terms, we can envision a trial state $\rho_{\mathrm{trial}}$ whose energy $\mathrm{tr}\!\left(H\rho_{\mathrm{trial}}\right)$ is of interest and supposed to approximate the ground state energy $\mathrm{tr}\!\left(H\rho_{\mathrm{gs}}\right)$. Then, we can use additional measurements to estimate $(1/Z)\mathrm{tr}\!\left(H\mathrm{e}^{-\tau H}\rho_{\mathrm{trial}}\mathrm{e}^{-\tau H}\right)$ with $Z=\mathrm{tr}\!\left(\mathrm{e}^{-2\tau H}\right)$, which is more representative of the ground state energy.
IV.4 Enhancing sub-universal quantum simulators
The underlying principle of this method is that both the generators of the time evolution operator and the observable belong to the same group. The limitation to short simulation times further means that we are restricted to simulating operators close to the identity. However, in contexts where one cannot perform arbitrary quantum operations, such as sub-universal quantum simulators, even operations close to the identity might be of interest, if they cannot be implemented natively. Therefore, as long as we can measure the generators of a unitary close to the identity and those generators belong to the same group as the observable of interest, we can apply this unitary in classical post-processing and thus expand the capabilities of the quantum simulator.
For example, imagine a quantum device that can only perform gates on neighboring qubits, but measure arbitrary single-qubit Pauli operators. Then, we can add long-range interactions in classical post-processing by, e.g., employing the presented method for $F(t)=\mathrm{e}^{\mathrm{i}H_{1}t_{1}}\cdots\mathrm{e}^{\mathrm{i}H_{r}t_{r}}O\mathrm{e}^{-\mathrm{i}H_{r}t_{r}}\cdots\mathrm{e}^{-\mathrm{i}H_{1}t_{1}}$, where the $H_{i}$ represent shallow sums of geometrically non-local Hermitian operators.
Together with the combination of imaginary time evolution with shortcuts to adiabaticity or device verification, these applications motivate classical post-processing techniques for analog or sub-universal quantum simulators. For such applications, we stress that focusing on Pauli operators is unnecessary and that, e.g., fermionic Hamiltonians and fermionic observables can be combined similarly.
V Discussion
The introduced Taylor series method to estimate short-time dynamics via measurements constitutes another feasible reading of the “measure now, compute functions of interest later” mindset, which combines quantum computations with classical post-processing of measurement statistics. Classical shadows once more present themselves as a multi-faceted tool that can help estimate several unrelated properties simultaneously enabling the search for new post-processing techniques, in this case allowing for a short-time quantum simulation through measurements application.
Although limited to short simulation times, it provides further arguments for the power of quantum measurements, adding to recent insights on the power of Pauli shadows on initial, untrained states in the context of quantum neural networks [49]. In this way, it is in line with other recent developments that aim at fleshing out the power of quantum measurement in quantum state engineering, quantum computation, and quantum simulation. It motivates the further development of hybrid quantum-classical algorithms and new ways to enhance them in post-processing. The limitation to short times is not unexpected since it is present in the purely classical regime of cluster expansion techniques and due to the classical hardness of time evolution in general.
The focus on Pauli operators is not required and decompositions of Hamiltonians into other operators following a group structure, such as Clifford operators, provides another path towards reducing resource costs and simplifying post-processing tasks. We hope the present work stimulates further rigorous research on quantum dynamics and hybrid quantum-classical algorithms.
Acknowledgments
We thank Mark Steudtner, Kianna Wan, Earl Campbell, Mario Berta, Jose Carrasco, and Alvaro M. Alhambra for useful discussions. The Berlin team has been supported by the BMWK (PlanQK), the DFG (CRC 183), the BMBF (QPIC-1, MuniQC-Atoms, HYBRID), the Quantum Flagship (Millenion, Pasquans2), QuantERA (HQCC) and the ERC (DebuQC). RiK has been supported by the Austrian Science fund via SFB BeyondC (Grant-DOI 10.55776/F71), the projects QuantumReady (FFG 896217) and High-Performance integrated Quantum Computing (FFG 897481) within Quantum Austria (both managed by the FFG), as well as the ERC (q-shadows).
Appendix A Tail bound for the truncated Taylor series
This appendix provides details of arguments presented in the main text. Using insights from the “Trotter error with 1-norm scaling” lemma of Ref. [39, Lemma 1], we can bound the operator norm of the remainder of the truncated Taylor series of the time-evolution operator by using Taylor’s remainder theorem with
$\left\lVert\sum_{k=K+1}^{\infty}\frac{(-\mathrm{i}Ht)^{k}}{k!}\right\rVert\leq\frac{\left(\lVert H\rVert t\right)^{K+1}}{(K+1)!},$ (19)
where we use the operator norm unless otherwise specified. Note that this is different from a naive tail bound that would result in a factor of $\mathrm{e}^{\lVert H\rVert t}$ instead. Furthermore, since
$(K+1)!\geq\frac{(K+1)^{K+1}}{\mathrm{e}^{K}}=\mathrm{e}\left(\frac{K+1}{\mathrm{e}}\right)^{K+1},$ (20)
we find
$\frac{\left(\lVert H\rVert t\right)^{K+1}}{(K+1)!}\leq\left(\frac{\lVert H\rVert t\mathrm{e}}{(K+1)}\right)^{K+1}\mathrm{e}^{-1}\leq\epsilon$ (21)
as soon as
$K+1=\frac{\log(1/(\mathrm{e}\epsilon))}{\mathrm{W}\left(\frac{1}{\lVert H\rVert t}\log(1/(\mathrm{e}\epsilon))\right)},$ (22)
where $\mathrm{W}$ denotes the Lambert $\mathrm{W}$ function defined as the inverse of $xe^{x}$. Assuming $\lVert Ht\rVert\leq 1$ yields a scaling of
$K=\mathcal{O}\left(\frac{\log(1/\epsilon)}{\log(\log(1/\epsilon)}\right),$ (23)
recovering the standard bound from Ref. [27, Eq. (4)]. When using $r$ segments of time evolution operators $U(t/r)$ with $r\geq\lVert H\rVert t$, the accuracy required for each segment is $\epsilon/r$, resulting in
$K=\mathcal{O}\left(\frac{\log(r/\epsilon)}{\log(\log(r/\epsilon)}\right).$ (24)
Appendix B Direct Taylor expansion of $O(t)$
In Section II.2, we have constructed an approximation of $O(t)$ by concatenating two truncated Taylor series of the time evolution operator. Consequently, the resulting approximation will also contain terms with higher powers of $t$ than the cutoff $K$. Although such an approximation will be more accurate, it differs from a direct Taylor expansion of the target function $F(H,t)=O(t)$. We will now showcase such a direct approach for the example of time dynamics and the resulting trade-off in accuracy and the number of terms in the linear combination of Pauli operators setting $r=1$ for simplicity.
B.1 Expansion
Instead of Eq. (4), we Taylor expand $O(t)$ directly as
$O(t)\!$ $=\!\sum_{k+k^{\prime}=0}^{K}\!\!\frac{\mathrm{i}^{k-k^{\prime}}t^{k+k^{\prime}}}{k!k^{\prime}!}H^{k}OH^{k^{\prime}}$
$\approx\sum_{k+k^{\prime}=0}^{K}\sum_{\stackrel{{\scriptstyle\bm{m}\in\llbracket 1..L_{H}\rrbracket^{k}}}{{\bm{n}\in\llbracket 1..L_{H}\rrbracket^{k^{\prime}}}}}\gamma_{\bm{m},\bm{n}}(t)Q_{\bm{m},\bm{n}},$ (25)
which again is a sum of $n$-qubit Pauli operators $Q_{\bm{m},\bm{n}}$ with corresponding coefficients $\gamma_{\bm{m},\bm{n}}(t)$, defined as
$\gamma_{\bm{m},\bm{n}}(t)$ $=\frac{\mathrm{i}^{k-k^{\prime}}t^{k+k^{\prime}}}{k!k^{\prime}!}\alpha_{m_{1}}\cdots\alpha_{m_{k}}\alpha_{n_{1}}\cdots\alpha_{n_{k^{\prime}}},$ (26)
$Q_{\bm{m},\bm{n}}$ $=H_{m_{1}}\cdots H_{m_{k}}OH_{n_{1}}\cdots H_{n_{k^{\prime}}}.$ (27)
The number of terms in this linear combination of Pauli operators is upper-bounded by $\sum_{k+k^{\prime}\leq K}L_{H}^{k+k^{\prime}}=\mathcal{O}(L^{K})$ and thus quadratically smaller than when concatenating two Taylor approximations. As such, we cannot expect the same accuracy of this approximation and must construct an appropriate bound on the systematic approximation error.
B.2 Error bound
We here present a rigorous error bound. To this end, we can use the “Trotter error with 1-norm scaling” lemma of Ref. [39]. This theorem bounds the tail $\mathrm{\textbf{R}}_{K}$ of product formulas above $K$-th order via
$\left\lVert\mathrm{\textbf{R}}_{K}\left(\prod_{\ell=1}^{L_{H}}\mathrm{e}^{\mathrm{i}\alpha_{\ell}H_{\ell}t}\right)\right\rVert$ $=\left\lVert\mathrm{e}^{\mathrm{i}Ht}-\prod_{\ell=1}^{L_{H}}\mathrm{e}^{\mathrm{i}\alpha_{\ell}H_{\ell}t}\right\rVert$ (28)
$=\frac{\left(\sum_{\ell=1}^{L_{H}}\left\lVert\alpha_{\ell}\,H_{\ell}\right\rVert t\right)^{K+1}}{(K+1)!}$
and can be adapted to find the error between the exact time evolution of some observable $O$ and the presented approximation, which can be bounded by
$\left\lVert\mathrm{e}^{-\mathrm{i}Ht}O\mathrm{e}^{\mathrm{i}Ht}-\widetilde{O}_{K}(t)\right\rVert$ $\leq\left\lVert\mathrm{\textbf{R}}_{K}\left(\mathrm{e}^{-\mathrm{i}Ht}O\mathrm{e}^{\mathrm{i}Ht}\right)\right\rVert$ (29)
$=\left\lVert\mathrm{\textbf{R}}_{K}\left(\mathrm{e}^{-\mathrm{i}Ht}\mathrm{e}^{\ln{O}}\mathrm{e}^{\mathrm{i}Ht}\right)\right\rVert$
$=\lVert O\rVert\frac{\left(2\lVert H\rVert t\right)^{K+1}}{(K+1)!},$
where we have lifted $O$ to fit the product formula form and recovered $\lVert O\rVert$ since Ref. [39, Lemma 6] has an additional factor of $\lVert\prod_{\ell=1}^{L_{H}}\mathrm{e}^{\mathrm{i}\alpha_{\ell}H_{\ell}t}\rVert$ which vanishes for all anti-Unitary operators in the exponent but remains for $\mathrm{e}^{\ln O}$. The error is thus larger by a factor $32^{k+1}/$ 3 compared to the concatenated approach discussed in Section II.3.1. Again, these results could be extended to a commutator bound using insights from Ref. [39].